Pure acid is to be added to a 10% acid solution to obtain 90 L of a 20% acid solution.
What amounts of each should be used?
____ L of pure acid solution and ____ L of 10% acid solution.

Answer by Alan3354(69443)

Pure acid is to be added to a 10% acid solution to obtain 90 L of a 20% acid solution.
What amounts of each should be used?
------------------
t = 10%
h = 100%
-----
t + h = 90 (total solution)
10t + 100h = 90*20 (total acid)
Solve for t & h